On August 1 of the current year, Kates employer Shop Inc. provided him a vehicle. For Shop Inc. to purchase the vehicle it would cost $45,000 including all applicable taxes. Whereas leasing the vehicle would cost Shop Inc. $750 a month. In either case, Shop Inc. would pay all operating costs for the vehicle, which is expected to be $2,250 yearly. Kates estimates that he will drive 2,100 monthly, of which 300 km will be for employment.

a) Determine the amount to be included in Kate's employment inform for the current year if i)

Shop Inc. purchases the car and ii) if Shop Inc. leases the car.

b) Kate drives the car for 16,000 km for employment purposes and 6,000 for personal use in the following year. Determine the amount to be included in his personal taxes assuming Shop Inc. leases the vehicle

Vehicle formulas

Operating expense benefit= prescribed rate (28 cents) X Personal KMS

Standby charge= Original cost of vehicle X 2% X # months available

Standby Charge = Monthly lease cost x 2/3 X 8 months available

Reduce Standby charge= (Personal km driven/1,667 km per month) X standby charge
Answer #1

Ans. a). Calculation of Amount to be included in Kate's Employment Form is given below-

(1) The Purchase Price of Car is $45000 and If taken on lease, the Annual Lease value is $750*12= $9000

(2) The Annual Operating Cost is $2250.

(3) Total Yearly Run KMs= 2100*12=25200 KMs. Out of this Yearly Run KM for Employment is 300*12= 3600 KM. So personal run of Kate is 21600 KMs (25200-3600).

Benefit to Kate, when car is purchased= 45000*21600/25200+ 2250*21600/25200= $40500. Under Vehicle Formula-

Operating Expense Benefit= 21600*28%= 6048 KMs. So $2250*6048/25200= $540

Standby Charge= $45000*2%*12= $ 10800.

So total benefit is $10800+$540= $11340.

Benefit to Kate, when car is taken on lease= $9000*21600/25200+2250*21600/25200= $9643.

Operating Expense Benefit= 21600*28%=6048 KMs. So $2250*6048/25200=$540.

Standby Charge= $750*2/3*8=$4000. So total benefit is $540+$4000=$4540.

Reduce Standby Charge= 300/1667*$4000= $720.

(b) Total Run KM= 16000+6000= 22000 KMs.

Total Yearly Cost= Total Yearly Lease Cost+ Annual Operating Cost= $750*12+$2250= $11250.

Benefit to Kate is $11250*6000/22000= $3068.

Operating Expense Benefit= 6000*28%= 1680 KMs. So Benefit is $11250*1680/22000= $859.

Standby Charge= $750*2/3*8=$4000

Reduce Standby Charge= 6000/1667*$4000= $14397.